a chemist has one solution that is 20% saline and a second that is 65% saline. How many gallons of each should be mixed together to get 120 gallons of a solution that is 50% saline.

Found 2 solutions by josgarithmetic, Alan3354:
Answer by josgarithmetic(39617)

and a second that is 65% saline.
?

Answer by Alan3354(69443)

a chemist has one solution that is 20% saline and a second that is 65% saline. How many gallons of each should be mixed together to get 120 gallons of a solution that is 50% saline.
=================
t = 20%
s = 65%
---
t + s = 120 (total solution)
20t + 65s = 120*50 (total salt)
---
Solve for t & s